Suppose income (Y) is given at 400 units, and the money supply (M) is fixed at 200 units.





Suppose k drops from its initial value of 0.5 to 0.25.





What is the initial price level? What is the new price level after the change in k?Explain the





process that leads to the change in the aggregate price level.




1
Expert's answer
2022-03-13T18:55:58-0400

"P1 = \frac{Md}{(k x Y)}=\frac{200}{0.5\times400}=1"


"P2 = \frac{Md}{(k x Y)}=\frac{200}{0.25\times400}=2"